Production Report

Dentex Plastics Corporation produces non-breakable container for cosmetics, using three departments; Mixing, Molding and Finishing.  On July 1, the work in process inventory in the Molding Department was 1,000 units, 50% complete as to materials and conversion costs, while the July 31 work in process inventory consisted of 2,800 units, 75% complete as to materials and conversion costs.  During, July, the Finishing Department received 20,000 units from the Molding Department.    Relevant cost data are as follows:

 

Work   in process July 1

Rs. 6,000

Cost added during July


Materials  

48,744

Labour

59,112

Factory   overhead

52,416

 

Required:

Using FIFO costing method and weighted average method, prepare the cost of production report for the Molding Department for July.
Answer #1
FIFO




Beginning1000
Goods inputted21800
Total accounted for22800



Completed and transferred20000
Ending2800
Total accounted for22800



FIFO


MaterialConversio
Beginning WIP10001000
% already complete50%50%
% to be completed50%50%
Equivalent beginning units500500
Units started and completed during the month1900019000
Ending WIP28002800
% to be completed75%75%
Equivalent ending units21002100
Total equivalent units2160021600



Total cost of equivalent units




Materials48744
Labor59112
Factory overhead52416
Total costs160272
Equivalent units21600
Cost per equivalent unit7.42



Cost of units completed and transferred out

Cost of equivalent unit144690=19500*7.42
Beginning WIP costs6000
Total costs150690



Cost of ending WIP15582=2100*7.42
Weighted average




Units completed and trasferred out2000020000
Ending WIP28002800
% to be completed75%75%
Equivalent ending units21002100
Total equivalent units2210022100



Cost per equivalent unit

Materials48744
Labor59112
Factory overhead52416
Cost of current month160272
Beginning costs6000
Total costs accounted for166272
Equivalent units22100
Cost per equivalent unit7.52362



Cost of units completed and transferred out

Cost of equivalent unit150472.4=20000*7.52362



Cost of ending WIP15799.6=2100*7.52362


answered by: studybay